Calculate relative rate of effusion of O2 to CH4 through a container containing O2 and CH4 in 3:2 mass ratio

Grade:11

1 Answers

Vikas TU
7 years ago
Rate of O2=  >
R1 = (3x/32)A/root(32T)
Rate of CH4 = >
R2 = (x/8)A/root(16T)
Divide both eqns.
we get,
R1/R2 = ¾ * root(16/32) = > 3/4root2 is the respective rate of effusion ratio.